Errata for Stable Marriage—first printing of the English edition (last updated 19 February 2015)

Page 5, bottom line a prefers A to b a prefers A to B Page 6, lines 17 and 18 the least preferred woman of V. the woman that V prefers least. Page 13, Point 6 stable stable. Page 18, lines 2 and 3 for Euler’s constant example for example Page 18, line 7 faster way fast way Page 20, lines 17 and 18 couple Cd. Since d prefers C to A, A becomes couple Cd, since d prefers C to A; A becomes Page 25, line 15 func˘ ion function

Page 28, line 5 from the bottom We find We find, using cumulative probabilities, Page 31, line 16 to city 1 from city 1 Page 33, line 2 of the third step citi˘ s cities

Page 39, line 13 from the bottom 10 ln 10/9 \frac{10}{9} ln 10 Page 41, line 1 Erdös Erdős Page 41, line 23 wheıı when Page 48, line 5 H, F, Q: array[1:n, 1:n] H, F, Q: array[1:n] Page 48, line 7 H[x] ← 0 H[X] ← 0 Page 52, line 8 if j = n if j = n Page 52, line 13 end end ; Page 52, line 23 F [X] ← X F [x] ← X Page 61, line 5 from the bottom the crossed-out c should occur first, the crossed-out e last (see page 57)

Page 61, line 3 from the bottom the circled e should be marked female-optimal, not male-optimal Page 63, line 7 x_{ij} = 1 ⇔ x_{ij} = 1 ⇒ Page 63, line 3 from the bottom q_{1}, … q_{n} q_{1}, …, q_{n} Page 73, right column Irving, Robert W. Irving, Robert Wylie